find moment of inertia of a circular ring about x-axis.
1
Expert's answer
2020-09-29T09:42:53-0400

Explanations & Calculations





  • Moment of inertia (I) is calculated by the relationship "I = \int r^2\delta m" about the axis in consideration.

When the x axis lies on the middle, (perpendicular to the plane of the ring) as the first figure,

  • Consider a small segment of mass "\small \delta m" of the ring then,

"\qquad\qquad
\begin{aligned}
\small I_1 &= \small \int r^2 \delta m\\
&= \small r^2\int_{0}^M \delta m\\
&= \small r^2 (M-0)\\&= \small Mr^2
\end{aligned}" : distance to the axis from each "\small \delta m" is the same ; r

When the x axis lies on a diameter of the ring consider a "\small \delta m_1" and linear mass density ("\small \rho = \frac{M}{2 \pi r}" ). Then,

"\qquad\qquad
\begin{aligned}
\small \delta m_1 &= \small \rho \delta l=\rho r\delta \theta\cdots (\because S=r\theta)\\

\small I_2 &= \small \int y^2 \delta m\\
&= \small \int (r\sin\theta)^2 \times \rho r\delta \theta \\
&= \small r^3\rho \int_{0}^{2\pi} \sin^2\theta \delta \theta\\
&= \small r^3\rho \int_{0}^{2\pi} \frac{1-\cos2\theta}{2} \delta \theta \\
&= \small \frac{r^3\rho}{2} \Bigg[\theta-\frac{\sin2 \theta}{2} \Bigg]_{0}^{2\pi}\\
&= \small \frac{r^3}{2}\times \frac{M}{2\pi r}\times 2\pi\\
&= \small \frac{Mr^2}{2}
\end{aligned}"